# {textblock*}

### v1

@@ 4.2 Implications for the Potential
The result (76) can be generalized to a wider class of potentials. As in the discussion around (55), it is useful to select an expression quantifying the steepness of the potential. In the present case, an appropriate choice is $-\partial_{\phi}V/V$. For potentials of the form (75), this is simply $\lambda$. It is thus intuitive (see also App. C.2) that for a given $V(\phi)$, brane-brane horizons exist if
@@
Furthermore, it is possible to consider potentials for which the decisive quantity $-\partial_{\phi}V/V$ asymptotes to $\lambda_{\rm crit}$ and study which of them induce brane-brane horizons. Concretely, it is shown in App. C.2 that potentials of the form
@@ 5 Conclusions
In this paper, we applied the Horizon Criterion, which identifies spacetimes with cosmological horizons as being problematic in quantum gravity, to spacetimes with boundaries. A spacetime $\mathcal{M}$ is said to have cosmological horizons if there exists an inertial observer that is causally disconnected from some parts of $\mathcal{M}$. If $\mathcal{M}$ has a dynamical boundary, i.e. an end-of-the-world (ETW) brane, such an observer can be bound to that brane. The most naive application of the Horizon Criterion would be to demand that all inertial observers, including both bulk and brane observers, do not experience a cosmological horizon. However, one quickly realizes that observers living on a Randall-Sundrum (RS) UV brane do see a cosmological horizon (cf. Fig. 1), as there are points in the AdS bulk spacetime that are causally disconnected from the ETW brane. Now, there is no fundamental reason to believe that the RS model is inconsistent in quantum gravity and, what is more, there exists an embedding in string theory [67]. It would be interesting to study the RS horizon from a brane-EFT or holographic point of view. In any case, the Horizon Criterion can not be applied in its most naive form.
@@
One possible conclusion from this observation would be that cosmological horizons simply do not represent a fundamental problem in quantum gravity. This takes us back to the originally discussed possibility that the problem is rooted in asymptotic acceleration. However, that appears unlikely since negatively curved universes accelerate even with steep potentials. Thus, maybe one should be more optimistic that de Sitter and similar cosmologies are not in the Swampland after all, or one must look for a better criterion excluding them.
@@
Alternatively, one would have to refine the Horizon Criterion. We argued that an appropriate generalization is to demand that any two observers living on the same ETW brane should be in causal contact with each other. To be more precise, we introduced the notion of a brane-brane horizon. By this we mean the intersection of the cosmological horizon of a brane-bound observer with the brane itself. The refined criterion then says that no brane-brane horizons should exist. We analyzed the consequences following from such a criterion for ETW branes whose action includes a brane localized potential $V(\phi)$, with $\phi$ a scalar field. For ETW branes bounding Anti-de Sitter (AdS) space, the corresponding conditions can be easily inferred using AdS/CFT correspondence. Hence, we focused on the more interesting case where a $(d\!-\!1)$-dimensional ETW brane bounds a $d$-dimensional geometry with vanishing cosmological constant. Brane-localized potentials with a positive energy minimum or approaching a positive value asymptotically trivially realize horizons and are hence excluded. Our main object of interest were hence brane-localized potentials $V(\phi)$ monotonically approaching zero from above at $\phi\to\infty$.
@@
It would be interesting to study string-theoretic realizations of our ETW brane setups, the stringy origin of non-zero potentials for a bulk scalar and their dynamical effects, and the consequences of dimensional reduction for the Horizon Criterion.
@@
By contrast, if horizons are indeed the property responsible for the difficulties with de Sitter, a better understanding of our brane-brane horizon criterion is needed: Ideally, one would like to identify more fundamental features distinguishing the RS from a dS horizon. One way forward may be analysing a ‘causal-isolation criterion’ that forbids cosmologies in which two inertial observers can be outside each other’s cosmological horizon in their asymptotic future. This criterion does not forbid the RS scenario: Here, the asymptotic future of the bulk observer lies outside the brane observer’s horizon, while the bulk observer does not even possess a horizon.
@@ A.1 Scalar Field living on the Brane
with $u_{0}$ an integration constant. We can now easily determine the profiles for $A(t),B(t)$. Going away from $z=0$ simply results in replacing $t$ with $t-z=u$. The corresponding profiles are:
@@ Appendix B Recovering the Potential from the Expansion Rate
In this section, we explain how to compute $V(\phi)$ if $H(\tau)$ or $H(R)$ is given. A similar analysis has been performed in [100, 101]. We restrict our attention to expanding universes, i.e. $H>0$, and to scalar field profiles that everywhere satisfy $\dot{\phi}>0$.
@@ Appendix C Autonomous System Analysis
In this appendix, we set $M_{d}=1$. For a review on the usage of autonomous systems in cosmology, see [74] and references therein.
@@ C.1 Scalar Field living on the Brane
[…]
@@
$\frac{dR}{d\tilde{\tau}}=\frac{\dot{R}}{H}=R\,,\qquad\Rightarrow\qquad R=\text{const}\cdot\exp(\tilde{\tau})\,.$ (121)
@@
[…]
@@
which can be seen as a measure for the local steepness of the potential. We note that for potentials of the form (117), $\delta=1/c$. By considering the variables $\delta$ and $U\equiv\dot{\phi}/\sqrt{2\rho}$, the dynamics of the system defined by (32) and (14) (together with (120)) can be studied via an autonomous system of equations taking the form (see e.g. [92, 96, 102, 103])
@@
If $f(\delta)$ vanishes at some $\delta=\delta_{*}$, there exists a fixed point given by [^13] In [104, 105], the function $f(\delta)$ was named $f$-deviser.
@@
The profile $\rho(\tilde{\tau})$ of the energy density, needed to evaluate the integral in (122), can be determined from the continuity equation (50), which takes the form
@@
A general result for autonomous systems, one of the center manifold theorems, see e.g. [106], implies that the solution for $y$ satisfies $y=\mathcal{O}(\kappa^{2})$. Therefore, it is safe to neglect the terms ($\order{\kappa^{2}y}$ ) in (135) and the asymptotic solutions for $\kappa$ and $u$ can be determined:
$y$ =($=\order{\tilde{\tau}^{-2}}\,,\qquad\kappa=-\frac{8\eta}{c_{\rm crit}\tilde{\tau}}+\mathcal{O}(\tilde{\tau}^{-2})\,,$ (136)
@@
[…]
@@ C.2 Scalar Field living in the Bulk
[…]
@@
$\zeta=\zeta_{*}\left(1-\frac{\beta}{\tilde{\tau}}\right)\,.$ (150)
@@ C.3 FLRW Universes
Using the definition (120), we find the autonomous system of equations (see e.g. [74, 107])
@@
$\frac{dy}{d\tilde{\tau}}$ $=-(d-2)y+h(\kappa)+\mathcal{O}\left(\kappa^{3},y^{2},\kappa y\right)\,,\qquad h(\kappa)=\order{\kappa^{2}}\,,$ (167)
@@
From the center-manifold theorems (see e.g. [106]), it follows that $y=\order{\kappa^{2}}$ ) such that the leading order solutions can be determined to be
$\kappa=-\frac{\beta}{\sqrt{d-2}\tilde{\tau}}+\order{\tilde{\tau}^{-2}}\,,\qquad u=-\frac{\beta}{2\sqrt{d-1}\tilde{\tau}}+\order{\tilde{\tau}^{-2}}\,.$ ) . (169)

### v2

@@ 4.2 Implications for the Potential
The result (76) can be generalized to a wider class of potentials. As in the discussion around (55), it is useful to select an expression quantifying the steepness of the potential. In the present case, an appropriate choice is $\partial_{\phi}V/V$. For potentials of the form (75), this is simply $\lambda$. It is thus intuitive (see also App. C.2) that for a given $V(\phi)$, brane-brane horizons exist if
@@
Furthermore, it is possible to consider potentials for which the decisive quantity $\partial_{\phi}V/V$ asymptotes to $\lambda_{\rm crit}$ and study which of them induce brane-brane horizons. Concretely, it is shown in App. C.2 that potentials of the form
@@ 5 Conclusions
In this paper, we extended the Horizon Criterion, which identifies spacetimes with cosmological horizons as being problematic in quantum gravity, to spacetimes with boundaries. A spacetime $\mathcal{M}$ is said to have cosmological horizons if there exists an inertial observer that is causally disconnected from some parts of $\mathcal{M}$. If $\mathcal{M}$ has a dynamical boundary, i.e. an end-of-the-world (ETW) brane, such an observer can be bound to that brane. The most naive extension of the Horizon Criterion would be to demand that all inertial observers, including both bulk and brane observers, do not experience a cosmological horizon. However, one quickly realizes that observers living on a Randall-Sundrum (RS) UV brane do see a cosmological horizon (cf. Fig. 1), as there are points in the AdS bulk spacetime that are causally disconnected from the ETW brane. Now, there is no fundamental reason to believe that the RS model is inconsistent in quantum gravity and, what is more, there exists an embedding in string theory Verlinde:1999fy. It would be interesting to study the RS horizon from a brane-EFT or holographic point of view. In any case, the Horizon Criterion can not be applied in its most naive form.
@@
One possible conclusion from this observation would be that cosmological horizons simply do not represent a fundamental problem in quantum gravity. One may also argue against horizon-type criteria because the definition of a horizon is intrinsically non-local. As a result, a certain amount of arbitrariness is introduced since one must decide to which set of observers it applies. This takes us back to the originally discussed possibility that the problem is rooted in asymptotic acceleration. However, that appears unlikely since negatively curved universes accelerate even with steep potentials. Thus, maybe one should be more optimistic that de Sitter and similar cosmologies are not in the Swampland after all, or one must look for a better criterion excluding them.
@@
Alternatively, one would have to refine the Horizon Criterion. We argued that an appropriate generalization is to demand that any two observers living on the same ETW brane should be in causal contact with each other. To be more precise, we introduced the notion of a brane-brane horizon. By this we mean the intersection of the cosmological horizon of a brane-bound observer with the brane itself. The Refined Horizon Criterion then says that no brane-brane horizons should exist. We analyzed the consequences following from such a criterion for ETW branes whose action includes a brane localized potential $V(\phi)$, with $\phi$ a scalar field. For ETW branes bounding Anti-de Sitter (AdS) space, the corresponding conditions can be easily inferred using AdS/CFT correspondence. Hence, we focused on the more interesting case where a (d1)(d\!-\!1)-dimensional ETW brane bounds a $d$-dimensional geometry with vanishing cosmological constant. Brane-localized potentials with a positive energy minimum or approaching a positive value asymptotically trivially realize horizons and are hence excluded. Our main object of interest were hence brane-localized potentials $V(\phi)$ monotonically approaching zero from above at $\phi\to\infty$.
@@
It would be interesting to study string-theoretic realizations of our ETW brane setups, the stringy origin of non-zero potentials for a bulk scalar and their dynamical effects, and the consequences of dimensional reduction for the Refined Horizon Criterion.
@@
By contrast, if horizons are indeed the property responsible for the difficulties with de Sitter, a better understanding of our Refined Horizon Criterion is needed: Ideally, one would like to identify more fundamental features distinguishing the RS from a dS horizon. One way forward may be analysing a ‘Causal-Isolation Criterion’ that forbids cosmologies in which two inertial observers can be outside each other’s cosmological horizon in their asymptotic future. This criterion does not forbid the RS scenario: Here, the asymptotic future of the bulk observer lies outside the brane observer’s horizon, while the bulk observer does not even possess a horizon. Another very interesting, even weaker version of a horizon-type criterion is the following: [^13] We thank the anonymous referee for pointing out this possibility.
One could demand that, among the observers with infinite worldline, at least one should have no horizon. The conjecture would then be that all solutions of a consistent theory of quantum gravity must satisfy this condition. Again, the RS solution would not be ruled out since bulk observers have no horizon.
@@ A.1 Scalar Field living on the Brane
with $u_{0}$ an integration constant. We can now easily determine the profiles for $A(t),B(t)$. Going away from $z=0$ simply results in replacing $t$ with tz=ut-z=u. The corresponding profiles are:
@@ Appendix B Recovering the Potential from the Expansion Rate
In this section, we explain how to compute $V(\phi)$ if $H(\tau)$ or $H(R)$ is given. A similar analysis has been performed in Starobinsky:1998fr, Copeland:2006wr. We restrict our attention to expanding universes, i.e. $H>0$, and to scalar field profiles that everywhere satisfy $\dot{\phi}>0$.
@@ Appendix C Autonomous System Analysis
In this appendix, we set $M_{d}=1$. For a review on the usage of autonomous systems in cosmology, see Bahamonde:2017ize and references therein.
@@ C.1 Scalar Field living on the Brane
$\infty>\int^{t_{\rm max}}dt=\int^{\infty}e^{-A(t)}d\tau=\int d\tau\frac{1}{e^{-A(t)}e^{B(t)}\partial_{t}B(t)}=\int^{\infty}\frac{d\tau}{\dot{R}(\tau)}\,$, (119)
@@
[…]
@@
$\infty>\int^{\infty}\frac{d\tau}{\dot{R}(\tau)}=\int^{\infty}\frac{d\tau}{HR}=\int^{\infty}\frac{d\tilde{\tau}}{H^{2}R}=\text{const}\cdot\int^{\infty}d\tilde{\tau}\frac{e^{-\tilde{\tau}}}{\rho^{2}}\,$, (122)
@@
which can be seen as a measure for the local steepness of the potential. We note that for potentials of the form (117), $\delta=1/c$. By considering the variables $\delta$ and $U\equiv\dot{\phi}/\sqrt{2\rho}$, the dynamics of the system defined by (32) and (14) (together with (120)) can be studied via an autonomous system of equations taking the form (see e.g. Mizuno:2002wa, Mizuno:2004xj, Goliath:1998na, Zhou:2007xp)
@@
If $f(\delta)$ vanishes at some $\delta=\delta_{*}$, there exists a fixed point given by [^15] In Escobar:2013js, Fadragas:2013ina, the function $f(\delta)$ was named $f$-deviser.
@@
[…]
@@
[…]
@@
$\infty>\int^{\infty}d\tilde{\tau}\tilde{\tau}^{-\frac{\eta c_{crit}^{2}}{d-2}}\,$, (139)
@@ C.2 Scalar Field living in the Bulk
$\infty>\int^{t_{\rm max}}dt$ $=\int^{\infty}e^{-A}d\tau=\int^{\infty}e^{-B}d\tau=\int^{\infty}\frac{d\tau}{R}=\int^{\infty}\frac{e^{-\tilde{\tau}}}{H}d\tilde{\tau}=\text{const.}\int^{\infty}\frac{e^{-\tilde{\tau}}}{V}d\tilde{\tau}\,$, (141)
@@
[…]
@@ C.3 FLRW Universes
Using the definition (120), we find the autonomous system of equations (see e.g. Bahamonde:2017ize, Copeland:1997et)
@@
[…]
@@
[…]